# Optimal or Greedy Decision Trees? Revisiting their Objectives, Tuning, and Performance

### v1

@@ 6 Conclusion
Future research could investigate the comparison of ODTs with greedy methods without explicit binarization as a preprocessing step. Additionally, the insights from the non-concave objectives could be further exploited, both for ODTs and greedy heuristic design.
Appendix A Data sets
@@
Table LABEL:tab:data_sets lists the OpenML 165 data sets used in the experiments in this paper (Vanschoren et al., 2013; Feurer et al., 2021).
@@
ID Data set Samples Features Binarized features
919 rabe_166 40 1 10
857 bolts 40 7 31
791 diabetes_numeric 43 2 20
729 visualizing_slope 44 3 30
928 rabe_97 46 3 13
777 sleuth_ex1714 47 7 70
817 diggle_table_a1 48 4 40
835 analcatdata_vehicle 48 4 9
942 chscase_health 50 3 27
787 witmer_census_1980 50 4 40
346 aids 50 4 30
476 analcatdata_bankruptcy 50 5 50
874 rabe_131 50 5 50
892 sleuth_case1201 50 6 60
780 rabe_265 51 6 60
713 vineyard 52 2 18
467 analcatdata_japansolvent 52 8 80
790 elusage 55 2 20
864 sleuth_ex2015 60 7 61
887 mbagrade 61 2 11
755 sleuth_ex1605 62 5 50
894 rabe_148 66 5 50
929 rabe_176 70 3 30
804 hutsof99_logis 70 7 37
1015 confidence 72 3 30
893 visualizing_hamster 73 5 50
859 analcatdata_gviolence 74 8 51
945 kidney 76 6 36
459 analcatdata_asbestos 83 3 14
472 lupus 87 3 26
946 visualizing_ethanol 88 2 14
40683 postoperative-patient-dat 88 8 22
1055 cm1_req 89 8 17
891 sleuth_case1202 93 6 44
731 baskball 96 4 40
865 analcatdata_neavote 100 2 10
875 analcatdata_chlamydia 100 3 16
726 fri_c2_100_5 100 5 50
1463 blogger 100 5 13
829 fri_c1_100_5 100 5 50
754 fri_c0_100_5 100 5 50
916 fri_c3_100_5 100 5 50
461 analcatdata_creditscore 100 6 37
45615 appendicitis_test_edsa 106 7 70
771 analcatdata_michiganacc 108 3 24
890 cloud 108 7 61
736 visualizing_environmental 111 3 30
782 rabe_266 120 2 19
448 analcatdata_boxing1 120 3 21
1556 acute-inflammations 120 6 15
714 fruitfly 125 4 24
40681 mux6 128 6 6
924 humandevel 130 1 10
867 visualizing_livestock 130 2 15
1075 datatrieve 130 8 71
885 transplant 131 3 30
921 analcatdata_seropositive 132 3 23
444 analcatdata_boxing2 132 3 21
974 hayes-roth 132 4 11
719 veteran 137 7 36
1013 analcatdata_challenger 138 2 12
784 newton_hema 140 3 30
902 sleuth_case2002 147 6 24
969 iris 150 4 39
955 tae 151 5 30
1026 grub-damage 155 8 48
40669 corral 160 6 6
748 analcatdata_wildcat 163 5 29
747 servo 167 4 19
801 chscase_funds 185 2 16
446 prnn_crabs 200 7 61
733 machine_cpu 209 6 49
796 cpu 209 7 58
895 chscase_geyser1 222 2 20
41538 conference_attendance 246 6 24
464 prnn_synth 250 2 20
730 fri_c1_250_5 250 5 50
744 fri_c3_250_5 250 5 50
911 fri_c2_250_5 250 5 50
776 fri_c0_250_5 250 5 50
1495 qualitative-bankruptcy 250 6 18
811 rmftsa_ctoarrivals 264 2 20
450 analcatdata_lawsuit 264 4 30
1073 jEdit_4.0_4.2 274 8 66
43 haberman 306 3 26
1524 vertebra-column 310 6 60
818 diggle_table_a2 310 8 79
1167 pc1_req 320 8 31
925 visualizing_galaxy 323 4 36
1011 ecoli 336 7 52
1048 jEdit_4.2_4.3 369 8 68
860 vinnie 380 2 13
1025 analcatdata_germangss 400 5 16
900 chscase_census6 400 6 60
906 chscase_census5 400 7 70
907 chscase_census4 400 7 70
908 chscase_census3 400 7 70
909 chscase_census2 400 7 70
1511 wholesale-customers 440 8 65
764 analcatdata_apnea3 450 3 20
814 chscase_vine2 468 2 18
724 analcatdata_vineyard 468 3 29
765 analcatdata_apnea2 475 3 20
767 analcatdata_apnea1 475 3 20
884 fri_c0_500_5 500 5 50
749 fri_c3_500_5 500 5 50
870 fri_c1_500_5 500 5 50
792 fri_c2_500_5 500 5 50
886 no2 500 7 70
750 pm10 500 7 70
335 monks-problems-3 554 6 15
333 monks-problems-1 556 6 15
949 arsenic-female-bladder 559 4 40
334 monks-problems-2 601 6 15
997 balance-scale 625 4 16
770 strikes 625 6 60
795 disclosure_x_tampered 662 3 30
931 disclosure_z 662 3 30
827 disclosure_x_noise 662 3 30
774 disclosure_x_bias 662 3 30
44596 blood-transfusion-service 748 4 34
44593 blood-transfusion-service 748 4 34
44594 blood-transfusion-service 748 4 34
44595 blood-transfusion-service 748 4 34
44597 blood-transfusion-service 748 4 34
1464 blood-transfusion-service 748 4 34
1014 analcatdata_dmft 797 4 20
912 fri_c2_1000_5 1000 5 50
813 fri_c3_1000_5 1000 5 50
743 fri_c1_1000_5 1000 5 50
799 fri_c0_1000_5 1000 5 50
45604 dummy 1000 6 60
43255 1StudentPerfromance 1000 7 43
741 rmftsa_sleepdata 1024 2 14
934 socmob 1156 5 31
1462 banknote-authentication 1372 4 40
991 car 1728 6 21
962 mfeat-morphological 2000 6 43
914 balloon 2001 1 10
772 quake 2178 3 30
40704 Titanic 2201 3 5
737 space_ga 3107 6 60
44127 phoneme 3172 5 49
871 pollen 3848 5 50
728 analcatdata_supreme 4052 7 24
720 abalone 4177 8 73
40983 wilt 4839 5 50
1460 banana 5300 2 20
803 delta_ailerons 7129 5 50
725 bank8FM 8192 8 73
816 puma8NH 8192 8 80
923 visualizing_soil 8641 4 31
819 delta_elevators 9517 6 47
44126 bank-marketing 10578 7 51
1046 mozilla4 15545 5 40
45558 Pulsar-Dataset-HTRU2 17898 8 80
45028 california 20634 8 80
45578 California-Housing-Classification 20640 8 80
843 house_8L 22784 8 75
45037 BitcoinHeist_Ransomware 24780 7 48
42493 airlines 26969 7 67
44156 electricity 38474 8 68
43901 click_prediction_small 39926 8 56
45022 Diabetes130US 71090 7 41
40922 Run_or_walk_information 88588 6 60

### v2

@@ 6 Conclusion
Future research could investigate the comparison of ODTs with greedy methods without explicit binarization as a preprocessing step.
Appendix A Splitting and Pruning Criteria as ODT Objectives
Greedy top-down induction (TDI) methods typically consider two types of criteria: a local splitting criterion that decides how to split a node and a pruning criterion that decides which nodes to keep during pruning. Some greedy methods employ another criterion for pruning such as the minimum description length (Mehta et al., 1995) or the pessimistic error rate in C4.5 (Quinlan, 1993). In this appendix, we list nine such criteria and rewrite them as ODT objectives.
Accuracy:
The most direct way to optimize out-of-sample accuracy is to optimize the in-sample accuracy by minimizing the misclassification score. The leaf node objective in terms of the leaf node size $n$ and the number of misclassifications $e$ is $f(n,e)=e$. A possible disadvantage of this objective is its ‘flatness’: it cannot distinguish between many different solutions. For example, $f(4,2)+f(6,0)=f(5,1)+f(5,1)$.
Gini impurity:
Commonly used in TDI heuristics is the Gini impurity. Weighted Gini impurity scores are obtained by multiplying the Gini impurity by the number of instances in that leaf node. Let $p_{0}=\frac{e}{n}$ denote the probability of the first class and $p_{1}=\frac{n-e}{n}$ the probability of the second class. Then the objective value is
$f(n,e)=n(1-p_{0}^{2}-p_{1}^{2})\,.$
Square root Gini:
$f(n,e)=n\sqrt{1-p_{0}^{2}-p_{1}^{2}}\,.$
Entropy:
The second commonly used TDI criterion is entropy (or information gain). Weighted scores are again obtained by multiplying by the size of the leaf node:
$f(n,e)=-\frac{n}{2}(p_{0}\log_{2}p_{0}+p_{1}\log_{2}p_{1})\,.$
Minimum error:
Niblett (1987) estimates the expected error for nodes by assuming that every class has equal probability. It depends on the number of labels $|\mathcal{K}|$, and the count of the majority label $n_{c}$. In binary classification $|\mathcal{K}|=2$ and $n_{c}=n-e$. Therefore, the expected error is
$n\frac{n-n_{c}+|\mathcal{K}|-1}{n+|\mathcal{K}|}=\frac{n(e+1)}{n+2}\,.$
Pessimistic error:
$f(n,e)=e+\frac{1}{2}\,.$
Binomial pessimistic error (Binom.):
In our experiments, we use the same default $\alpha=0.25$ as C4.5.
Minimum description length (Quinlan):
$L(n,e,b)=\ln{(b+1)}+\ln{\binom{n}{e}}\,.$ (2)
$f(n,e)=L(n,e,\lfloor(n+1)/2\rfloor)\,.$
Minimum description length (Mehta):
Bayesian:
$f(n,e)=\frac{B(e+\rho_{0},n-e+\rho_{1})}{B(\rho_{0},\rho_{1})}\,.$
Almost all previous ODT methods optimize the accuracy, but some have considered other objectives, such as balanced accuracy (Lin et al., 2020) or F1-score (Demirović and Stuckey, 2021; van der Linden et al., 2023). Nijssen and Fromont (2010) have implemented the pessimistic error objective from C4.5 and a Bayesian objective function, but do not discuss their effect on the out-of-sample performance.
Appendix B ODT Tuning Approaches
This appendix provides further details on the ODT hyperparameter tuning approaches, the choice of hyperparameters when limited to $k$ options, and more experiments.
B.1 Details on Existing Tuning Approaches
We evaluate four existing ODT hyperparameter tuning approaches: depth, size, cost-complexity, and minimum support. Here, we provide some background on each of the tuning methods, and how we select the $k$ options for the hyperparameter. For each approach, we always include a setting that provides no limitation on the tree size. The other settings are derived using linear or log-equal distances between settings. The following provides more details per method:
Depth:
A common metric for tree complexity is its depth. DL8.5 (Aglin et al., 2020a, b), for example, tunes the depth of the tree. A possible disadvantage is that the number of nodes increases exponentially with respect to the maximum depth, thus providing only a coarse control of the tree size.
In our experiments, we tune the depth parameter $d\in\{0,...,\operatorname{max-depth}\}$ with equal linear space between the options and with a $\operatorname{max-depth}$ of either four or five.
Size:
The number of nodes can be tuned directly as a hard constraint. Since binary trees always have one more leaf node than branching nodes, the total number of nodes can be counted by the number of branching or leaf nodes alone. This approach is taken by, e.g., MurTree (Demirović et al., 2022) which directly tunes both the maximum depth and the number of branching nodes.
[…]
Complexity cost:
The most common approach in greedy trees is to penalize the cost of adding a node by a factor $\lambda$. This approach is used in most MIP methods (Bertsimas and Dunn, 2017) and also in the optimal sparse approaches such as GOSDT (Lin et al., 2020). Complexity-cost tuning minimizes $\lambda|\mathcal{D}||L|+\sum_{(n,e)\in L}f(n,e)$, with $L$ the set of leaves, $|\mathcal{D}|$ the size of the data set, and $\lambda$ the complexity-cost parameter.
The minimum change in $\lambda$ that may result in a different tree is $\lambda=\frac{1}{|\mathcal{D}|\operatorname{max-depth}}$ because below this value it is never worth adding nodes to increase accuracy. We set the maximum value to $0.05$ and select $k-1$ options from this range using equal log spacing. The minimum step size between values is set to $\frac{1}{|\mathcal{D}|\operatorname{max-depth}}$. In all cases, we add $\lambda=0$.
Note that Lin et al. (2020) recommend for their method GOSDT to use $\lambda\geq\frac{1}{|\mathcal{D}|}$ for faster training, but this setting can exclude larger trees that are more accurate. Additionally, in their experiments, they aim to acquire trees of at most $n$ leaves by setting $\lambda=\frac{1}{n}$. However, this also filters out many trees that have much less than $n$ leaves, since generally a single leaf node already has an accuracy greater than 0.5 for binary classification, so that even a perfect tree with $n$ leaves given such $\lambda$ would have a lower score than a single leaf node. Therefore, both of these settings are too conservative.
Minimum support:
The minimum support is a hard constraint on the minimum number of instances that in a leaf node. Minimum support is more common in the data mining literature, e.g., Nijssen and Fromont (2007). It is also introduced as a soft constraint (Vilas Boas et al., 2021).
In our experiments, we tune the minimum leaf node size $m$ as a percentage of the data set size $|\mathcal{D}|$. The minimum value for $m$ is $\frac{1}{|\mathcal{D}|}$: precisely one sample should end up in each leaf node. For the maximum value, we compute the frequency of the majority class $|\mathcal{D}_{\operatorname{majority}}|$ and set the maximum value of $m$ to $1-\frac{|\mathcal{D}_{\operatorname{majority}}|}{|\mathcal{D}|}$. Any value higher than this will always result in a single leaf node. We obtain $k$ values from this range using equal log spacing, with a minimum step size of $\frac{1}{|\mathcal{D}|}$.
We leave out selecting a tree structure from a fixed set (Günlük et al., 2021), because this approach was chosen specifically to deal with the scalability limitations of MIP.
B.2 Details on the New Tuning Approaches
We set up the two new tuning approaches introduced in the main text as follows:
Question length:
In our experiments, we select $k-1$ options for $\omega$ using equal log spacing from the interval $[\frac{1}{|\mathcal{D}|\operatorname{max-depth}},0.1]$. Additionally, we always add the option $\omega=0$. We also use $\frac{1}{|\mathcal{D}|\operatorname{max-depth}}$ as the minimum step size.
Smoothing:
In our experiments, we select $k-1$ options for $x$ using equal log spacing from the interval $[\frac{1}{\operatorname{max-depth}},0.05|\mathcal{D}|]$. Additionally, we always add the option $x=0$. We also use $\frac{1}{\operatorname{max-depth}}$ as the minimum step size.
B.3 Extended Tuning Experiments
Here, we report additional experiments on the ODT tuning methods. We test the performance of the methods for other values of $k$, the number of hyperparameter options, and we report the average runtime for each approach. While in the main text, we experimented with $k=16$, here we test for $k=5,10,20$. For runtime comparison, we ran all experiments on an Intel Xeon E5-6248R 3.0GHz with 100GB RAM using one thread.
Figure 20: Performance of all tuning methods for $k=5,10,20$ with $\operatorname{max-depth}=4$ on the 180 OpenML data sets. The color indicates the tuning method. Test accuracy is roughly the same for all methods. Increasing $k$ typically yields smaller trees. As expected, runtime scales roughly linear with $k$.
Fig. 20 shows that the choice of $k$ has no major impact on the out-of-sample performance of the tuning methods. In almost all cases, running with more hyperparameter options results in a lower average number of leaf nodes. As expected, increasing $k$ by a factor two, also increases the runtime for each method by a factor of roughly two.
These results also show that the runtimes of all methods are in the same order of magnitude. The only exception to this is tuning the depth. Therefore, the choice of hyperparameter tuning and the number of parameter options to test mostly depends on how much time one is willing to spend on finding small trees and less important for optimizing accuracy.
Appendix C Previous Comparisons between Optimal and CART
C.1 Comparisons in ODT Papers
C.2 Other Comparisons
Murthy and Salzberg (1995) compare the greedy approach to known true (optimal) trees on a synthetic benchmark. They observe that the greedy tree is approximately one standard deviation larger than the true tree size while the question length (what they call expected depth) is very close to optimal, which was also observed by Goodman and Smyth (1988) (Claim 6). However, the maximum depth of greedy trees is on average approximately two times higher than the true depth. When they increase the data set size linearly with the true tree complexity, they observe almost no drop in accuracy for the greedy approach (Claim 4). From this result, Costa and Pedreira (2023) hypothesize that the gap between optimal and greedy approaches diminishes for more data (Claim 3).
Appendix D Additional Experiments
Section D.1 provides more details on the train and test accuracy of various objectives. Section D.2 measures the effect of binarization for ODTs.
D.1 Train Accuracy
Figure 21: Train and test accuracy for ODTs with different objectives for increasing size limits for the data sets visualizing_slope ($n$=44), aids ($n$=50), confidence ($n$=72), and lupus ($n$=87). Train accuracy is always close. Test accuracy differences are larger.
Figure 22: Train and test accuracy for ODTs with different objectives for increasing size limits for the data sets servo ($n$=167), haberman ($n$=306), vinnie ($n$=380), and pollen ($n$=3848). For haberman, vinnie, and pollen, the train accuracy of Gini impurity is worse.
Fig. 21 and 22 show the train and test accuracies obtained by training optimal decision trees for three objectives for an increasing node limit for small and medium-sized data sets. All trees are trained with $\operatorname{max-depth}=4$. The training accuracy of the three objectives is typically close. For the larger data sets, Gini impurity obtains a lower training accuracy. Optimizing accuracy or M-loss results in (almost) the same training accuracy.
For several of these data sets, the training accuracy plateaus early. For some (other) data sets, the test accuracy remains stable or decreases when trained with more nodes, indicating that the trees are possibly overfitting. The variance in the test accuracy results is large, making it hard to draw conclusions from these figures.
D.2 Binarization
To test the effect of binarizing features, we experiment with changing the number of binary features per original numeric or categorical feature. For all data sets, we compute optimal trees of maximum depth three or four when creating 2, 5, 10, 15, or 25 binary features per original feature. We then measure the test accuracy for each tree and compute the normalized test accuracy by dividing it by the highest test accuracy obtained for that data set. Fig. 23 shows how on average the test accuracy plateaus when creating more and more binary features per original feature. Our selection of 10 binary features per original feature on average is 0.6% less accurate for $\operatorname{max-depth}=3$ and 0.5% less accurate for $\operatorname{max-depth}=4$ than when using 15 binary features per original feature. Fig. 23 also shows the exponential increase in runtime when the number of binary features is increased.
Figure 23: The normalized accuracy for an increasing number of binary features per original feature. At some point, more binary features add little extra information, while significantly increasing the runtime.
What binarization technique works best is a topic that we leave aside in this paper. Other work studies this, such as McTavish et al. (2022) who use ensembles for binarization, and Piccialli et al. (2024) who use counterfactuals for discretization. Recently, also new ODT methods have been proposed that do not require explicit binarization (Mazumder et al., 2022; Shati et al., 2023; Brita et al., 2025).
In this paper, the precise method of binarization is not a major concern. We assume that the binarization is given, so that all methods work on precisely the same data. In this way, we eliminate this difference between methods and can focus in the comparison on the difference between the greedy and optimal approaches.
Appendix E Data sets
@@
Table LABEL:tab:data_sets lists the 180 OpenML data sets used in the experiments in this paper (Vanschoren et al., 2013; Feurer et al., 2021).
@@
ID Data set Samples Features Binarized features Class imbalance
919 rabe_166 40 1 10 0.53
857 bolts 40 7 31 0.65
938 sleuth_ex1221 42 9 90 0.55
40660 analcatdata_fraud 42 11 19 0.69
791 diabetes_numeric 43 2 20 0.60
729 visualizing_slope 44 3 30 0.61
777 sleuth_ex1714 47 7 70 0.57
835 analcatdata_vehicle 48 4 9 0.56
817 diggle_table_a1 48 4 40 0.52
942 chscase_health 50 3 27 0.52
346 aids 50 4 30 0.50
787 witmer_census_1980 50 4 40 0.52
476 analcatdata_bankruptcy 50 5 50 0.50
892 sleuth_case1201 50 6 60 0.52
713 vineyard 52 2 18 0.54
467 analcatdata_japansolvent 52 8 80 0.52
790 elusage 55 2 20 0.56
864 sleuth_ex2015 60 7 61 0.55
887 mbagrade 61 2 11 0.52
755 sleuth_ex1605 62 5 50 0.50
804 hutsof99_logis 70 7 37 0.51
1015 confidence 72 3 30 0.83
893 visualizing_hamster 73 5 50 0.55
859 analcatdata_gviolence 74 8 51 0.58
945 kidney 76 6 36 0.53
459 analcatdata_asbestos 83 3 14 0.55
472 lupus 87 3 26 0.60
862 sleuth_ex2016 87 10 80 0.52
946 visualizing_ethanol 88 2 14 0.51
40683 postoperative-patient-dat 88 8 22 0.73
1055 cm1_req 89 8 17 0.78
479 analcatdata_cyyoung9302 92 9 67 0.79
891 sleuth_case1202 93 6 44 0.61
731 baskball 96 4 40 0.51
465 analcatdata_cyyoung8092 97 10 78 0.75
865 analcatdata_neavote 100 2 10 0.93
875 analcatdata_chlamydia 100 3 16 0.81
1463 blogger 100 5 13 0.68
754 fri_c0_100_5 100 5 50 0.54
46544 dataset_analcatdata_credi 100 6 36 0.73
965 zoo 101 16 20 0.59
45615 appendicitis_test_edsa 106 7 70 0.80
771 analcatdata_michiganacc 108 3 24 0.56
890 cloud 108 7 61 0.70
736 visualizing_environmental 111 3 30 0.52
448 analcatdata_boxing1 120 3 21 0.65
1556 acute-inflammations 120 6 15 0.51
714 fruitfly 125 4 24 0.61
40681 mux6 128 6 6 0.50
924 humandevel 130 1 10 0.50
867 visualizing_livestock 130 2 15 0.81
1075 datatrieve 130 8 71 0.92
885 transplant 131 3 30 0.63
921 analcatdata_seropositive 132 3 23 0.65
974 hayes-roth 132 4 11 0.61
719 veteran 137 7 36 0.69
1013 analcatdata_challenger 138 2 12 0.93
784 newton_hema 140 3 30 0.50
902 sleuth_case2002 147 6 24 0.53
1006 lymph 148 18 47 0.55
969 iris 150 4 39 0.67
955 tae 151 5 30 0.66
1026 grub-damage 155 8 48 0.68
40669 corral 160 6 6 0.56
748 analcatdata_wildcat 163 5 29 0.71
747 servo 167 4 19 0.77
463 backache 180 31 99 0.86
801 chscase_funds 185 2 16 0.53
941 lowbwt 189 9 37 0.52
1012 flags 194 28 117 0.64
446 prnn_crabs 200 7 61 0.50
733 machine_cpu 209 6 49 0.73
796 cpu 209 7 58 0.75
895 chscase_geyser1 222 2 20 0.60
41538 conference_attendance 246 6 24 0.87
464 prnn_synth 250 2 20 0.50
776 fri_c0_250_5 250 5 50 0.50
1495 qualitative-bankruptcy 250 6 18 0.57
811 rmftsa_ctoarrivals 264 2 20 0.62
450 analcatdata_lawsuit 264 4 30 0.93
336 SPECT 267 22 22 0.79
1073 jEdit_4.0_4.2 274 8 66 0.51
23499 breast-cancer-dropped-mis 277 9 37 0.71
1121 badges2 294 10 43 0.71
40710 cleve 303 13 70 0.54
43 haberman 306 3 26 0.74
1524 vertebra-column 310 6 60 0.68
818 diggle_table_a2 310 8 79 0.53
1167 pc1_req 320 8 31 0.67
925 visualizing_galaxy 323 4 36 0.54
1011 ecoli 336 7 52 0.57
1048 jEdit_4.2_4.3 369 8 68 0.55
860 vinnie 380 2 13 0.51
1025 analcatdata_germangss 400 5 16 0.78
909 chscase_census2 400 7 70 0.51
1511 wholesale-customers 440 8 65 0.68
1498 sa-heart 462 9 79 0.65
814 chscase_vine2 468 2 18 0.55
724 analcatdata_vineyard 468 3 29 0.56
4329 thoracic_surgery 470 16 54 0.85
767 analcatdata_apnea1 475 3 20 0.87
884 fri_c0_500_5 500 5 50 0.50
750 pm10 500 7 70 0.51
886 no2 500 7 70 0.50
40690 threeOf9 512 9 9 0.54
335 monks-problems-3 554 6 15 0.52
333 monks-problems-1 556 6 15 0.50
949 arsenic-female-bladder 559 4 40 0.86
826 sensory 576 11 32 0.59
334 monks-problems-2 601 6 15 0.66
997 balance-scale 625 4 16 0.54
770 strikes 625 6 60 0.50
774 disclosure_x_bias 662 3 30 0.52
827 disclosure_x_noise 662 3 30 0.50
795 disclosure_x_tampered 662 3 30 0.51
931 disclosure_z 662 3 30 0.53
40981 Australian 690 14 80 0.56
1464 blood-transfusion-service 748 4 34 0.76
37 diabetes 768 8 73 0.65
1014 analcatdata_dmft 797 4 20 0.81
44268 anneal 898 38 104 0.54
50 tic-tac-toe 958 9 27 0.65
40693 xd6 973 9 9 0.67
799 fri_c0_1000_5 1000 5 50 0.50
45604 dummy 1000 6 60 0.73
43255 1StudentPerfromance 1000 7 43 0.52
741 rmftsa_sleepdata 1024 2 14 0.50
40702 solar-flare 1066 10 26 0.83
40706 parity5_plus_5 1124 10 10 0.50
934 socmob 1156 5 31 0.78
40680 mofn-3-7-10 1324 10 10 0.78
1462 banknote-authentication 1372 4 40 0.56
983 cmc 1473 9 35 0.57
40646 GAMETES_Epistasis_2-Way_2 1600 20 60 0.50
40649 GAMETES_Heterogeneity_20a 1600 20 58 0.50
991 car 1728 6 21 0.70
962 mfeat-morphological 2000 6 43 0.90
914 balloon 2001 1 10 0.76
772 quake 2178 3 30 0.56
40704 Titanic 2201 3 5 0.68
737 space_ga 3107 6 60 0.50
44127 phoneme 3172 5 49 0.50
3 kr-vs-kp 3196 36 38 0.52
871 pollen 3848 5 50 0.50
728 analcatdata_supreme 4052 7 24 0.76
720 abalone 4177 8 73 0.50
40983 wilt 4839 5 50 0.95
45039 compas-two-years 4966 11 34 0.50
44160 rl 4970 12 69 0.50
1460 banana 5300 2 20 0.55
803 delta_ailerons 7129 5 50 0.53
43922 mushroom 8124 22 109 0.52
807 kin8nm 8192 8 80 0.51
725 bank8FM 8192 8 73 0.60
816 puma8NH 8192 8 80 0.50
923 visualizing_soil 8641 4 31 0.55
819 delta_elevators 9517 6 47 0.50
44126 bank-marketing 10578 7 51 0.50
4534 PhishingWebsites 11055 30 46 0.56
45060 online_shoppers 12330 17 136 0.85
959 nursery 12960 8 26 0.67
1046 mozilla4 15545 5 40 0.67
44162 compass 16644 17 111 0.50
45558 Pulsar-Dataset-HTRU2 17898 8 80 0.91
45028 california 20634 8 80 0.50
823 houses 20640 8 80 0.57
43904 law-school-admission-bian 20800 10 53 0.68
843 house_8L 22784 8 75 0.70
45037 BitcoinHeist_Ransomware 24780 7 48 0.50
42493 airlines 26969 7 67 0.55
43900 amazon_employee_access 32769 9 90 0.94
44156 electricity 38474 8 68 0.50
44120 electricity 38474 7 61 0.50
137 BNG(tic-tac-toe) 39366 9 27 0.65
43901 click_prediction_small 39926 8 56 0.83
881 mv 40768 10 75 0.60
46554 Loan_Status 45000 13 94 0.78
45547 Cardiovascular-Disease-da 70000 11 49 0.50
45022 Diabetes130US 71090 7 41 0.50
40922 Run_or_walk_information 88588 6 60 0.50
Figure 24: Histogram of the number of instances and binarized features in the data sets considered in our experiments.
@@
Fig. 24 shows a histogram of the number of instances and the number of binarized features of the data sets considered in this paper.